OpenStudy (anonymous):

cos314degrees 40' = a. -.7030 b. .7112 c. .7030 d. -.7112

OpenStudy (anonymous):

1st, its in the 4th quadrant so cos is +ve right?

OpenStudy (anonymous):

@mathsux4real now convert the 40' to degrees: 60' =1degree 40' = ??

OpenStudy (ash2326):

We have \[\cos (314\ degrees, 40')\] we know \[\cos (270+x)=\sin x\] so it's \[\sin (44\ degrees, 40')\] we know that \[\sin 45=\frac{1}{\sqrt 2}=0.707\] so \[\sin (44\ degrees, 40')<0.707\ ( \text{since 44 degrees, 40'<45 degrees})\] I think you can find the answer now

OpenStudy (anonymous):

Quadrant IV so cos is +, 40' in degrees is .666, now cos of 314.66(degrees) is .703

OpenStudy (anonymous):

@mathsux4real 40' = 40/60degree = 2/3 remember its in 4th quad so ref angle = 360- 314 2 /3 = 45 1/3 = 45.333 cos 45.3333= 0.7030

OpenStudy (ash2326):

@mathsux4real we don't need calculator for this, did you follow my explanation?

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

Wish I could give more than one medal